OpenStudy (anonymous):

PLEASE HELP ME Mandy's Candies is most famous for their homemade fudge and homemade toffee. They make an average of 150 pounds of fudge and toffee combined each month. They sell a pound of fudge for $8.84 and a pound of toffee for $7.63. Mandy's averages $1,231.62 each month in fudge and toffee sales. On average, how many pounds each of fudge and toffee does Mandy's Candies sell in one month? 78 fudge; 72 toffee 75 fudge; 75 toffee 72 fudge; 78 toffee 61 fudge; 89 toffee

OpenStudy (anonymous):

@acxbox22

OpenStudy (jhannybean):

Solve this as a system of linear equations. \[\begin{align} x=\text{toffee}~,~ y=\text{fudge} \\ x+y = 150 &\\ 8.84x+7.63y = 1231.62&\\ \end{align}\]

OpenStudy (jhannybean):

So you can use the elimination method to solve for one type of candy, and then use that to find the second type of candy.

OpenStudy (anonymous):

ok D?

OpenStudy (jhannybean):

\[-8.84(x+y=150)\]\[8.84x+7.63y=1231.62\] \[-8.84y +7.63y =-94.38\]\[-1.21y=-94.38\]\[~~~~~~~~~~y=78\]

OpenStudy (anonymous):

Since it can be almost any number of pounds, use the method of elimination on the given answers. You know the total pounds are supposed to be 150, and the total price is 1231.62 Multiply each of these by their corresponding price, and check the results. Example for the first one: 78 x $8.84 + 72 x $7.63 = $1238.88

OpenStudy (jhannybean):

So with y = 78, plug that in to the first equation to find x. x + 78 = 150 x = 150 - 78 = ?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

@Jhannybean

OpenStudy (jhannybean):

Yep.

OpenStudy (anonymous):

thank you